# Cristian Sestito  Shady Agwa  and Themis Prodromakis

### v1

@@ Abstract
Modern hardware architectures for Convolutional Neural Networks (CNNs), other than targeting high performance, aim at dissipating limited energy. Reducing the data movement cost between the computing cores and the memory is a way to mitigate the energy consumption. Systolic arrays are suitable architectures to achieve this objective: they use multiple processing elements that communicate each other to maximize data utilization, based on proper dataflows like the weight stationary and row stationary. Motivated by this, we have proposed TrIM, an innovative dataflow based on a triangular movement of inputs, and capable to reduce the number of memory accesses by one order of magnitude when compared to state-of-the-art systolic arrays. In this paper, we present a TrIM-based hardware architecture for CNNs. As a showcase, the accelerator is implemented onto a Field Programmable Gate Array (FPGA) to execute the VGG-16 CNN. The architecture achieves a peak throughput of 453.6 Giga Operations per Second, outperforming a state-of-the-art row stationary systolic array by 5.1$\sim$ 5.1$\times$ 5.1 × in terms of memory accesses, and being up to 12.2$\sim$ 12.2$\times$ 12.2 × more energy-efficient than other FPGA accelerators.
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
In this work, we propose an AI hardware architecture based on TrIM, dealing with multi-dimensional CLs for CNNs. As a case study, the TrIM architecture is implemented onto a Field Programmable Gate Array (FPGA) to accelerate the VGG-16 CNN at 150 MHz clock frequency. The design, consisting of 1512 PEs, exhibits a peak throughput of 453.6 GOPs/s, by dissipating $\sim$4.2 W of power. In addition, the TrIM architecture outperforms Eyeriss[23] by 5.1$\sim$ 5.1$\times$ 5.1 × in terms of memory accesses, and shows the best energy efficiency among state-of-the-art FPGA counterparts.
@@
  • A high-level scalable TrIM-based AI hardware architecture is presented. It incorporates 7 cores, each consisting of 24 slices of $3\times 3$ PEs.
  • A reconfigurable shift-register buffer is presented. It handles different ifmap sizes for optimal SA utilization. This unlocks the TrIM-based architecture to cope with all the CLs of a generic CNN, being run-time reconfigured.
@@
  • An FPGA-based implementation is presented to characterize the TrIM architecture in terms of resource occupation, performance, power and energy. Comparisons with Eyeriss and state-of-the-art FPGA architectures are also presented, in order to highlight the benefits offered by the TrIM-based architecture.
@@ II Related Works
In the spectrum of hardware-based SAs, the TrIM-based architecture sits at the GeMM-free side, by proposing an empowered WS dataflow where inputs are reused among different PEs through an innovative triangular dataflow. This allows a significant reduction of memory accesses. Multiple instances of $K\times K$ PEs are used, with $K$ being the kernel size of the 3-D convolutional filters that accommodate the weights. Few extra shift register buffers assist the inputs reuse.
@@ III The TrIM Architecture
The TrIM-based AI hardware architecture is designed as a hierarchy consisting of three levels:
@@
  • TrIM Engine. This top-level module consists of $P_{N}$ cores, each dealing with an independent 3-D filter. $P_{N}$ is the number of filters processed in parallel and, in turn, the number of ofmaps generated in parallel. All the cores use the same set of ifmaps.
@@ III-A TrIM Slice
The TrIM slice includes an array of $K\times K$ PEs, an adder tree and $K-1$ Reconfigurable Shift Register Buffers (RSRBs), as illustrated in Fig. 3. The computing array is organized in $K$ rows ($Row_{i}$, with $0\leq i<K$), each having $K$ PEs. Preliminarily, weights are fetched from the memory and provided to the PEs of $Row_{0}$ as groups of $K$ elements per cycle, then shifted from top to down. This process lasts $K$ cycles to ensure that the current kernel is completely stored at the slice level. Thus, inputs are first grabbed from the periphery ($I_{ext}$), moved in each row from right ($I_{R}$) to left ($I_{L}$), and forwarded to the closest SRB to complete the triangular movement. Each SRB dispatches $K$ inputs ($I_{D}$) to $Row_{i-1}$. Psums are first accumulated vertically through PEs and, finally, through an adder tree that finalizes the $K\times K$ convolution.
@@
The Adder Tree is supplied by $K$ psums coming from the bottom row of the array for the final accumulation. It consists of $\lceil\log_{2}(K)\rceil$ stages, with the last one followed by an output register.
@@ III-B TrIM Core
The subsequent adder tree produces a $2\times B+K+\lceil\log_{2}(K)\rceil+\lceil\log_{2}(P_{M})\rceil$-bit result (core_out), followed by an output register. Although the adder tree complexity can constraint the speed, pipelining can be added for shorter critical paths and better timing closure.
@@ IV Case Study: VGG-16
Figure 7: Design Space Analysis: (a) Throughput and psums buffers size considering the parallelism factors $P_{N}$ and $P_{M}$. Five groups of bars refer to as many values of $P_{N}=[1,4,8,16,24]$ (i.e., parallel filters/ofmaps). For each group, five different bars are related to as many values of $P_{M}=[1,4,8,16,24]$ (i.e., parallel ifmaps), as reported into the legend. Throughput is expressed in terms of tera operations per second. The dotted line indicates the psums buffers size, which is only dependent by $P_{N}$; (b) I/O bandwidth considering the parallelism factos $P_{N}$ and $P_{M}$, in Mbits. The bars follow the same format of the plot (a).
@@ V Hardware Implementation
TABLE I: TrIM vs Eyeriss: Performance Comparison on VGG-16
@@
CL No. $H_{O}\times W_{O}$ $M$ $N$ GOPs/s Memory Accesses [M] PE Utilization GOPs/sa Memory Accesses [M]b PE Utilization
1 $224\times 224$ 3 64 51.8 4.53 0.13 13.7 18.77 0.93
2 $224\times 224$ 64 64 368 41.92 1.00 13.7 400.47 0.93
3 $112\times 112$ 64 128 387 20.29 1.00 13.7 200.23 0.93
4 $112\times 112$ 128 128 387 40.87 1.00 13.7 400.47 0.93
5 $56\times 56$ 128 256 396 20.60 1.00 27.2 101.23 0.93
6 $56\times 56$ 256 256 432 40.42 1.00 27.2 202.47 0.93
7 $56\times 56$ 256 256 432 40.42 1.00 27.2 202.47 0.93
8 $28\times 28$ 256 512 422 21.91 1.00 52.8 53.63 1.00
9 $28\times 28$ 512 512 422 43.87 1.00 52.8 107.28 1.00
10 $28\times 28$ 512 512 422 43.87 1.00 52.8 107.28 1.00
11 $14\times 14$ 512 512 389 13.34 1.00 57.4 15.00 1.00
12 $14\times 14$ 512 512 389 13.34 1.00 57.2 15.00 1.00
13 $14\times 14$ 512 512 389 13.34 1.00 57.2 15.00 1.00
 Total 391 358.71 0.93 24.5 1839.30 0.94
aRetrieved from the number of VGG-16 operations per layer and following the given processing latency.
bRetrieved from the global memory accesses, and scaled on one image and 16-bit accesses.
The TrIM Engine has been synthesized and implemented onto the FPGA of the AMD Zynq UltraScale+ XCZU7EV-2FFVC1156 MultiProcessor System-on-Chip (MPSoC). The design has been carried out through the AMD Vivado 2022.2 tool, and using Verilog as Hardware Description Language.
First, taking into account the design space reported in Section IV, the optimal parameters $P_{M}$ and $P_{N}$ have been retrieved. $P_{N}$ is constrained by the on-chip memory. Considering that the XCZU7EV part offers 11 Mb of Block RAMs (BRAMs), using equation (3), we got $P_{N}=7$, supposing that each psums buffer stores up to 224 $\times$ 224 output activations, which refer to the worst-case scenario (first two layers of VGG-16). The parameter $P_{M}$ strictly depends on the available I/O bandwidth. The XCZU7EV device is interfaced to a 64-bit DDR4 memory and exhibits a peak bandwidth of 19200 MB/s. Supposing $f_{CLK}$ = 150 MHz, in line with state-of-the-art FPGA-based accelerators[25, 21, 24], $BW_{I/O}$ = 1024, rounded to the closest power of 2. Therefore, considering equation (4) and $P_{N}=7$, we got $P_{M}=24$. Hence, a TrIM Engine hosting $P_{N}=7$ cores, each having $T_{M}=24$ slices, has been synthesized and implemented onto FPGA. The TrIM-based architecture achieves a peak throughput of 453.6 GOPs/s, since 1512 PEs are responsible for performing $7\times 24$ parallel convolutions. With specific reference to the VGG-16 model, the engine guarantees a throughput of 391 GOPs/s, only 13.8% lower than the peak throughput. This is motivated by high PE utilization, which reaches the 93% on average.
In order to highlight the advantages offered by TrIM in terms of performance and local data utilization, we present a comparative analysis with Eyeriss[23], state-of-the-art systolic array proposing the RS dataflow. Table I details such a comparison, be reporting (a) the specific configuration of each CL, (b) throughput, memory accesses, and utilization for both the proposed architecture (TrIM) and Eyeriss. Eyeriss, despite working at 1.3×1.3$\times$1.3 × higher frequency, reaches a throughput 16$\sim$ 16$\times$ 16 × lower, because using only 168 PEs.
TABLE II: State-of-The-Art FPGA Architectures Based on Systolic Arrays
@@
Power [W] 11 1.398 8.25 4.219
Energy Efficiency [GOPs/s/W] 37.24 54.94 8.84 107.51
@@
TrIM requires 358.71 millions of memory accesses to retrieve and send ifmaps, weights, psums, ofmaps. Specifically, ifmaps account for 259.26 millions of accesses, weights for 14.03 millions, psums for 72.5 millions and 12.92 millions for ofmaps. On average, TrIM outperforms Eyeriss by 5.1$\sim$ 5.1$\times$ 5.1 ×, thanks to the effectiveness of the triangular movement of inputs at the slice level. Indeed, the inputs overhead per slice (i.e., the number of extra accesses) is only the $\sim 5.8\%$ of the total memory accesses. Furthermore, it worth to mention that Eyeriss exhibits an additional contribution related to the scratch pads, sitting at the PE level. Considering the consumption breakdown reported in [23], it has been estimated that scratch pads contribute from 12.9$\sim$ 12.9$\times$ 12.9 × to 16.5$\sim$ 16.5$\times$ 16.5 × more than the external memory.
@@
The TrIM-based architecture is also compared to some state-of-the-art accelerators implemented onto FPGA [25, 21, 24]. Table II summarizes the results, considering: (a) device type, data precision, number of processing elements and dataflow; (b) area occupation, in terms of Look-Up Tables (LUTs), Flip-Flops (FFs), Digital Signal Processing slices (DSPs), and Block RAMs (BRAMs); (c) clock frequency, peak throughput in GOPs/s; (d) power and energy efficiency.
@@
Sense[25] exploits a SA consisting of 32$\times$32 PEs. Considering that each PE must ensure both the WS and OS dataflow, other than guaranteeing the management of sparse convolutions, this results in higher hardware complexity and, in turn, more area. Indeed, even though TrIM makes use of $\sim$1.5$\times$ more PEs, Sense requires $\sim$79.9and $\sim$72.9more LUTs and BRAMs, respectively. In addition, while TrIM completely manages multiply-accumulations through cheap logic offered by LUTs, Sense adopts DSPs. In turn, this affects the power consumption and, thus, the energy efficiency, being 3$\sim$ 3$\times$ 3 × lower than TrIM.
@@
The architecture presented in[21] proposes a SA using 8$\times$32 PEs and dealing with the WS dataflow. Among the counterparts, this hardware engine is the cheapest in terms of logic and on-chip memory, except for using 257 DSPs to implement the PEs. In turn, this results as the lowest power footprint. However, considering that [21] reaches a peak throughput 5.9×5.9$\times$5.9 × lower than TrIM, this reflects in an energy efficiency 2$\sim$ 2$\times$ 2 × lower.
@@
The SA introduced in[24] makes use of 3$\times$3$\times$27 PEs using the RS dataflow. In particular, it proposes a scheduling module that arranges the rows of data beforehand, thus simplifying the hardware complexity of PEs with respect to Eyeriss. However, despite using $\sim$6.2$\times$ less PEs than TrIM, this scales only $\sim$2$\times$ in terms of LUTs, because of extra logic requirements to manage sparse convolutions. Overall, TrIM outperforms [24] by 12.2×12.2$\times$12.2 × in terms of energy efficiency.

### v2

# TrIM, Triangular Input Movement Systolic Array for Convolutional Neural Networks:

@@ Abstract
Modern hardware architectures for Convolutional Neural Networks (CNNs), other than targeting high performance, aim at dissipating limited energy. Reducing the data movement cost between the computing cores and the memory is a way to mitigate the energy consumption. Systolic arrays are suitable architectures to achieve this objective: they use multiple processing elements that communicate each other to maximize data utilization, based on proper dataflows like the weight stationary and row stationary. Motivated by this, we have proposed TrIM, an innovative dataflow based on a triangular movement of inputs, and capable to reduce the number of memory accesses by one order of magnitude when compared to state-of-the-art systolic arrays. In this paper, we present a TrIM-based hardware architecture for CNNs. As a showcase, the accelerator is implemented onto a Field Programmable Gate Array (FPGA) to execute the VGG-16 and AlexNet CNNs. The architecture achieves a peak throughput of 453.6 Giga Operations per Second, outperforming a state-of-the-art row stationary systolic array up to 3$\sim$ 3$\times$ 3 × in terms of memory accesses, and being up to 11.9$\sim$ 11.9$\times$ 11.9 × more energy-efficient than other FPGA accelerators.
@@ I Introduction
In this work, we propose an AI hardware architecture based on TrIM, dealing with multi-dimensional CLs for CNNs. As a case study, the TrIM architecture is implemented onto a Field Programmable Gate Array (FPGA) to accelerate the VGG-16 and AlexNet CNNs at 150 MHz clock frequency. The design, consisting of 1512 PEs, exhibits a peak throughput of 453.6 GOPs/s, by dissipating $\sim$4.3 W of power. In addition, the TrIM architecture outperforms Eyeriss[23] by up to 3$\sim$ 3$\times$ 3 × in terms of memory accesses, and shows the best energy efficiency among state-of-the-art FPGA counterparts.
@@
  • A high-level scalable TrIM-based AI hardware architecture is presented. This includes innovative TrIM Slices that implement the dataflow by a hardware perspective. In order to execute CLs, these slices are arranged in 7 cores, each consisting of 24 slices of $3\times 3$ PEs. The cores compress psums spatially and the memory bandwidth is fully utilized by reading inputs once and broadcasting them to the different cores. Finally, different slices may cooperate with each other to manage large kernel sizes.
  • A reconfigurable shift-register buffer is presented. It handles different ifmap sizes for optimal SA utilization. This unlocks the TrIM-based architecture to cope with all the CLs of a generic CNN, being run-time reconfigured. Additionally, it is worth pointing out that the use of shift-register buffers and the data broadcasting at the top level allow the architecture to manage the memory bottleneck at different hierarchical levels.
@@
  • An FPGA-based implementation is presented to characterize the TrIM architecture in terms of resource occupation, performance, power and energy. VGG-16 and AlexNet CNNs are considered, allowing validation on different ifmap and kernel sizes. Comparisons with Eyeriss and state-of-the-art FPGA architectures are also presented, in order to highlight the benefits offered by the TrIM-based architecture.
@@ II Related Works
In the spectrum of hardware-based SAs, the TrIM-based architecture sits at the GeMM-free side. The TrIM dataflow combines IS and WS dataflows to maximize local data utilization, without affecting the external memory capacity. The TrIM dataflow operates on $K\times K$ PEs, interconnected with each other. Weights, inputs and psums go through the array as follows:
  • Weights are grabbed from the memory and progressively stored at the PE level from top to bottom. The stored weights are thus kept stationary for the entire convolution.
  • Inputs are initially read from the memory and moved into the array through vertical links. Given that convolutional sliding windows share elements column-wise and row-wise, inputs are locally reused through horizontal movements or diagonal movements, respectively. Horizontal reuse is achieved from the rightmost PE to the leftmost PE, while diagonal reuse is supported by $K-1$ Shift Register Buffers. These buffers provisionally store row-wise shared inputs between sliding windows. These data are thus supplied diagonally to PEs. The composition of the three movements (i.e., vertical, horizontal, diagonal) configures a triangular shape, thus the name TrIM.
  • Psums are accumulated from top to bottom. The psums generated at the bottom row of the array are finally reduced through an adder tree.
The local reuse of inputs make TrIM effective to mitigate the memory access count. For example, a $3\times 3$ convolution over a $224\times 224$ ifmap exhibits a negligible 1.8% overhead. In the following Section, we introduce a hardware architecture dealing with TrIM and suitable for the acceleration of CLs.
@@ III The TrIM Architecture
The TrIM-based AI hardware architecture is designed as a hierarchy consisting of three levels, as reported in Fig. 2. The modules composing the hierarchy are the following:
@@
  • TrIM Engine. This top-level module consists of $P_{N}$ cores, each dealing with an independent 3-D filter. $P_{N}$ is the number of filters processed in parallel and, in turn, the number of ofmaps generated in parallel. All cores use the same set of ifmaps.
@@ III-A TrIM Slice
The TrIM slice includes an array of $K\times K$ PEs, an adder tree and $K-1$ Reconfigurable Shift Register Buffers (RSRBs), as illustrated in Fig. 3. The computing array is organized in $K$ rows ($Row_{i}$, with $0\leq i<K$), each having $K$ PEs. Preliminarily, the weights are taken from memory and provided to the PEs of $Row_{0}$ as groups of $K$ elements per cycle, then shifted from top to bottom. This process lasts $K$ cycles to ensure that the current kernel is completely stored at the slice level. Thus, inputs are first grabbed from the periphery ($I_{ext}$), moved in each row from right ($I_{R}$) to left ($I_{L}$), and forwarded to the closest SRB to complete the triangular movement. Each SRB dispatches $K$ inputs ($I_{D}$) to $Row_{i-1}$. Psums are first accumulated vertically through PEs and, finally, through an adder tree that finalizes the $K\times K$ convolution.
@@
The Adder Tree is supplied by $K$ psums coming from the bottom row of the array for the final accumulation. Based on the binary-tree architecture, it consists of $\lceil\log_{2}(K)\rceil$ stages, with the last one followed by an output register.
@@ III-B TrIM Core
The subsequent adder tree, based on the binary-tree architecture, produces a $2\times B+K+\lceil\log_{2}(K)\rceil+\lceil\log_{2}(P_{M})\rceil$-bit result (core_out), followed by an output register. Although the adder tree complexity can constraint the speed, pipelining can be added for shorter critical paths and better timing closure.
@@ IV Case Study: VGG-16
Figure 7: Design Space Analysis: (a) Throughput and psums buffers size considering the parallelism factors $P_{N}$ and $P_{M}$. Five groups of bars refer to as many values of $P_{N}=[1,4,8,16,24]$ (i.e., parallel filters/ofmaps). For each group, five different bars are related to as many values of $P_{M}=[1,4,8,16,24]$ (i.e., parallel ifmaps), as reported into the legend. Throughput is expressed in terms of tera operations per second. The points indicate the psums buffers size, which is only dependent by $P_{N}$; (b) I/O bandwidth considering the parallelism factos $P_{N}$ and $P_{M}$, in Mbits. The bars follow the same format of the plot (a).
@@ V Hardware Implementation
TABLE I: TrIM vs Eyeriss: Comparison on VGG-16
@@
Memory Accesses [M]a Memory Accesses [M]a
CL No. $H_{I}\times W_{I}$ $K$ $M$ $N$ GOPs/s PE Util. On-Chipb Off-Chip Total GOPs/sc PE Util. On-Chipb Off-Chip Total
1 $224\times 224$ 3 3 64 51.8 0.13 0.00 13.57 13.57 13.7 0.93 43.81 7.70 51.51
2 $224\times 224$ 3 64 64 368 1.00 0.57 102.79 103.36 13.7 0.93 477.14 27.00 504.14
3 $112\times 112$ 3 64 128 387 1.00 0.27 49.96 50.23 13.7 0.93 271.44 16.70 288.14
4 $112\times 112$ 3 128 128 387 1.00 0.68 95.33 96.01 13.7 0.93 495.48 24.25 519.73
5 $56\times 56$ 3 128 256 396 1.00 0.33 48.51 48.84 27.2 0.93 145.57 10.10 155.67
6 $56\times 56$ 3 256 256 432 1.00 0.66 94.71 95.38 27.2 0.93 259.22 16.10 275.32
7 $56\times 56$ 3 256 256 432 1.00 0.66 94.71 95.38 27.2 0.93 255.46 15.40 270.86
8 $28\times 28$ 3 256 512 422 1.00 0.33 52.44 52.77 52.8 1.00 89.08 8.90 97.98
9 $28\times 28$ 3 512 512 422 1.00 0.70 103.72 104.42 52.8 1.00 157.88 14.30 172.18
10 $28\times 28$ 3 512 512 422 1.00 0.70 103.72 104.42 52.8 1.00 141.23 11.40 152.63
11 $14\times 14$ 3 512 512 389 1.00 0.17 33.05 33.23 57.4 1.00 32.69 3.15 35.84
12 $14\times 14$ 3 512 512 389 1.00 0.17 33.05 33.23 57.2 1.00 29.68 2.85 32.53
13 $14\times 14$ 3 512 512 389 1.00 0.17 33.05 33.23 57.2 1.00 28.95 2.80 31.75
 Total 391 0.93 5.44 858.63 864.06 24.5 0.94 2427.63 160.65 2588.28
aNormalized to the same data width and a batch of 3 images.
bNormalized to off-chip memory accesses.
cRetrieved from the number of VGG-16 operations per layer and following the given processing latency.
TABLE II: TrIM vs Eyeriss: Comparison on AlexNet
TrIM Eyeriss[23]
Memory Accesses [M]a Memory Accesses [M]a
CL No. $H_{I}\times W_{I}$ $K$ $M$ $N$ GOPs/s PE Util. On-Chipb Off-Chip Total GOPs/sc PE Util. On-Chipb Off-Chip Total
1 $227\times 227$ 11 3 96 2.13 1.00 0.08 8.44 8.52 51.1 0.92 17.92 2.50 20.42
2 $27\times 27$ 5 48 256 179 0.57 0.21 3.50 3.71 45.7 0.80 28.64 2.00 30.64
3 $13\times 13$ 3 256 384 390 1.00 0.11 14.85 14.95 54.9 0.93 15.09 1.50 16.59
4 $13\times 13$ 3 192 384 402 1.00 0.07 11.20 11.27 56.1 0.93 10.44 1.05 11.49
5 $13\times 13$ 3 192 256 399 1.00 0.05 7.52 7.57 59.8 0.93 5.36 0.65 6.01
 Total 12.9 0.91 0.53 45.50 46.03 51.5 0.88 77.45 7.70 85.15
aNormalized to the same data width and a batch of 4 images.
bNormalized to off-chip memory accesses.
cRetrieved from the number of AlexNet operations per layer and following the given processing latency.
TABLE III: State-of-The-Art FPGA Architectures for Systolic Arrays
@@
Power [W] 11 1.398 8.25 4.329
Energy Efficiency [GOPs/s/W] 37.24 54.94 8.84 104.78
@@
The TrIM-based architecture has been synthesized and implemented onto the FPGA of the AMD Zynq UltraScale+ XCZU7EV-2FFVC1156 MultiProcessor System-on-Chip (MPSoC). The design has been carried out through the AMD Vivado 2022.2 tool, and using Verilog as Hardware Description Language.
First, taking into account the design space reported in Section IV, the optimal parameters $P_{M}$ and $P_{N}$ have been retrieved. $P_{N}$ is constrained by the on-chip memory. Considering that the XCZU7EV part offers 11 Mb of Block RAMs (BRAMs), using equation (3), we got $P_{N}=7$, supposing that each psums buffer stores up to 224 $\times$ 224 output activations, which refer to the worst-case scenario (first two layers of VGG-16). The parameter $P_{M}$ strictly depends on the available I/O bandwidth. The XCZU7EV device is interfaced to a 64-bit DDR4 memory and exhibits a peak bandwidth of 19200 MB/s. Supposing $f_{CLK}$ = 150 MHz, in line with state-of-the-art FPGA-based accelerators[25, 21, 24], $BW_{I/O}$ = 1024, rounded to the closest power of 2. Therefore, considering equation (4) and $P_{N}=7$, we got $P_{M}=24$. Hence, a TrIM Engine hosting $P_{N}=7$ cores, each having $P_{M}=24$ slices, has been synthesized and implemented onto FPGA. Placement and routing has been fully managed by Vivado. The architecture uses 194.35K Look-Up Tables (LUTs), 89.72K Flip-Flops (FFs) and 10.21 Mb of on-chip memory through Block RAMs (BRAMs). The pipeline depth of the TrIM Engine is 9 stages, of which: 5 stages for the TrIM Slices, 3 stages for the adder tree at the core level, and 1 stage for the temporal accumulation at the engine stage. The dynamic power is 4.329 W (86% computations and data movement; 4% on-chip memory; 10% clock). The TrIM-based architecture achieves a peak throughput of 453.6 GOPs/s, since 1512 PEs are responsible for performing $7\times 24$ parallel convolutions. With specific reference to the VGG-16 model, the engine guarantees a throughput of 391 GOPs/s, only 13.8% lower than the peak throughput. This is motivated by high PE utilization, which reaches the 93% on average.
In order to highlight the advantages offered by TrIM in terms of performance and local data utilization, we present a comparative analysis with Eyeriss[23], state-of-the-art systolic array proposing the RS dataflow. To showcase the effectiveness of TrIM in a wide spectrum of ifmap sizes and kernel sizes, the VGG-16[7] and the AlexNet[29] CNNs are benchmarked. Table I and II detail such a comparison, reporting (a) the specific configuration of each CL, (b) throughput, PE utilization and memory accesses. Regarding the memory accesses, a breakdown between on-chip and off-chip memory is reported, and normalized to off-chip memory accesses.
Regarding VGG-16, TrIM takes 78.6 ms (391 GOPs/s) to perform one inference step, while Eyeriss takes 1.25 s (24.5 GOPs/s). In terms of total memory accesses, TrIM requires 3$\sim$ 3$\times$ 3 × less than Eyeriss. This mitigation is motivated by the very limited contribution of on-chip memory. In fact, while Eyeriss requires a global buffer and scratch pads at the PE level, TrIM only uses a global buffer for psums. In detail, the $\sim 94\%$ of equivalent on-chip memory accesses relates to scratch pads in the Eyeriss architecture. On the other hand, the use of a global buffer to store ifmap and weight tiles, along with data compression for ifmap sparsity, allows Eyeriss to save 5.3×5.3$\times$5.3 × off-chip memory accesses than TrIM. However, this is not sufficient to hide the on-chip memory contribution, which is 15$\sim$ 15$\times$ 15 × higher.
To cope with the different kernel sizes required by AlexNet, the TrIM architecture splits large kernels in $3\times 3$ tiles. For example, $P_{M}$ $5\times 5$ kernels are split in 4 groups of $P_{M}$ tiles each. Each group is processed by a TrIM Core and the psums are accumulated at the top level to produce the ofmap associated to the entire $5\times 5$ kernel. TrIM takes 103.1 ms to peform one inference step, while Eyeriss takes 26 ms. In terms of throughput, Eyeriss outperforms TrIM only in the first CL due to a mitigation of parallelism to split $11\times 1$ 11 kernels in $3\times 3$ tiles. However, in the rest of layers ($5\times 5$ and $3\times 3$ kernels) TrIM outperforms Eyeriss up to 7×7$\times$7 ×. The drawback related to the first layer can be mitigated by more hardware specialization for the large kernels. In terms of memory accesses, TrIM confirms the trend achieved with the VGG-16 CNN. In the AlexNet case, TrIM uses 1.8$\sim$ 1.8$\times$ 1.8 × less memory accesses than Eyeriss. In fact, the equivalent on-chip memory accesses (scratch pads + global buffer) are 10$\sim$ 10$\times$ 10 × higher than off-chip accesses in Eyeriss. Also in this case, the motivation sits at the Eyeriss’ PE level, where scratch pads dominate the on-chip accesses.
@@
The TrIM-based architecture is also compared to some state-of-the-art accelerators implemented onto FPGA [25, 21, 24]. Table III summarizes the results, considering: (a) device type, data precision, number of processing elements and dataflow; (b) area occupation, in terms of LUTs, FFs, Digital Signal Processing slices (DSPs), and BRAMs; (c) clock frequency, peak throughput in GOPs/s; (d) power and energy efficiency.
@@
Sense[25] exploits a SA consisting of 32$\times$32 PEs. Considering that each PE must ensure both the WS and OS dataflow, other than guaranteeing the management of sparse convolutions, this results in higher hardware complexity and, in turn, more area. Indeed, even though TrIM makes use of $\sim$1.5$\times$ more PEs, Sense requires $\sim$79.1more LUTs. In addition, while TrIM completely manages multiply-accumulations through cheap logic offered by LUTs, Sense adopts DSPs. TrIM, indeed, uses 8 bits for inputs and weights. Thus, computations do not require expensive 48-bit DSPs. In turn, this affects the power consumption and, thus, the energy efficiency, being 3$\sim$ 3$\times$ 3 × lower than TrIM.
@@
The architecture presented in[21] proposes a SA using 8$\times$32 PEs and dealing with the WS dataflow. Among the counterparts, this hardware engine is the cheapest in terms of logic and on-chip memory, except for using 257 DSPs to implement the PEs. In turn, this results as the lowest power footprint. However, considering that [21] reaches a peak throughput 5.9×5.9$\times$5.9 × lower than TrIM, this reflects in an energy efficiency 1.9$\sim$ 1.9$\times$ 1.9 × lower.
@@
The SA introduced in[24] makes use of 3$\times$3$\times$27 PEs using the RS dataflow. In particular, it proposes a scheduling module that arranges the rows of data beforehand, thus simplifying the hardware complexity of PEs with respect to Eyeriss. However, despite using $\sim$6.2$\times$ less PEs than TrIM, this scales only $\sim$1.8$\times$ in terms of LUTs, because of extra logic requirements to manage sparse convolutions. Overall, TrIM outperforms [24] by 11.9×11.9$\times$11.9 × in terms of energy efficiency.